# IterResearch: Rethinking Long-Horizon Agents via Markovian State Reconstruction

### v1

@@ 3.2.1 Discounted Reward Shaping for Efficiency
While the Markovian workspace reconstruction ensures scalable exploration, a critical question remains: how can we train agents to not just explore deeply, but to do so efficiently? We now address this challenge by introducing an efficiency-aware policy optimization framework.
In deep research tasks, the agent receives a binary reward signal $R_{T}\in\{0,1\}$ only upon termination, where $R_{T}=1$ if the final answer is correct and 0 otherwise. This terminal-only reward stems from the inherent difficulty of evaluating intermediate research steps—it is challenging to determine the value of any particular search query or exploratory action [chen2025cpo].
However, this sparse signal alone is insufficient for guiding efficient learning, as it treats all successful trajectories equally regardless of their computational cost. An agent that arrives at the correct answer in 5 well-chosen steps should be preferred over one that requires 20 steps of meandering exploration, even if both ultimately succeed. This efficiency consideration is not merely about computational resources: in real-world deployment, each interaction incurs API costs, and unnecessary exploration can lead to increased latency. To address these issues, we introduce a reward shaping mechanism based on geometric discounting from MDP theory [bellman1957markovian]:
@@ 3.2.2 Policy Optimization with Multi-Round Trajectories
This approach ensures minimal data loss (typically $1\%$ of samples) while maintaining uniform sampling across trajectories. To optimize IterResearch, we integrate our geometric discounted rewards and adaptive downsampling with the Group Sequence Policy Optimization (GSPO) algorithm [zheng2025group], enabling stable training on variable-length trajectories:
@@
where all $\sum_{i=1}^{G}T_{i}$ rounds from the $G$ trajectories for question $q$ form one group, with normalized advantages computed across all samples within this group $\hat{A}_{i,t}=\frac{r_{i,t}-\mu_{r}}{\sigma_{r}}$, $\mathcal{Q}$ is the training set, and $\rho_{i,t}(\theta)$ is the importance ratio based on sequence likelihood [zheng2023click].
@@ 4.1 Experimental Setup
Datasets. To rigorously assess the effectiveness of our IterResearch, we evaluate on six challenging benchmarks including Humanity’s Last Exam (HLE) [hle], BrowseComp [bc_en], BrowseComp-zh [bc_zh], GAIA [mialon2023gaia], Xbench-DeepSearch [xbench], SEAL-0 [pham2025sealqa]. These benchmarks comprehensively assess the essential capabilities for effective deep research in multi-step tool use, web navigation, complex reasoning, long-horizon information-seeking, and cross-lingual synthesis.
Baselines. We comprehensively compare our IterResearch against state-of-the-art methods including: (1) Direct Inference: We evaluate frontier LLMs including GPT-4o and GPT-4.1 [gpt_4o], o4-mini [openai_o3_o4_mini], and DeepSeek-R1-0528 [guo2025deepseek]. (2) Proprietary Deep-Research System: We compare with commercial deep-research systems including OpenAI’s Deep Research [dr], Perplexity Research [Perplexity], Gemini Deep Research [google_dr], Grok3-ResearchSearch [grok3], and Kimi-Researcher [kimi-researcher]. (3) Open-source Agents: Recent open-source deep-research agents including Search-o1 [li2025search], WebThinker [Li2025webthinker], WebDancer [wu2025webdancer], WebSailor [li2025websailor], Asearcher [gao2025beyond], and MiroThinker [MiroThinker].
Implementation Details. We implement our IterResearch using Qwen3-30B-A3B [yang2025qwen3] as the backbone model, considering both model performance and computational efficiency. Our training follows a two-stage process: we first employ rejection sampling fine-tuning (RFT) [yuan2023scaling] to equip the model with our iterative deep-research paradigm capabilities, then apply reinforcement learning to further enhance its search strategy and reasoning abilities. For brevity, we provide comprehensive training details and hyperparameters in Appendix C.3.
@@ 4.4 Scaling on Interaction
First, performance scales gracefully with interaction budget. Accuracy improves from 5.5% with only 2 turns to 50.1% at 2048 turns, with the steepest gains occurring between $2^{4}$ and $2^{7}$ turns. This demonstrates that complex information-seeking tasks genuinely benefit from extended exploration—a capability that mono-contextual approaches cannot provide due to inevitable context overflow. Notably, 2048 turns represents an extreme challenge that is currently infeasible for mono-contextual agents due to catastrophic context accumulation, yet our approach operates smoothly within its constant 40K token workspace through Markovian state reconstruction. Second, the agent learns intelligent resource allocation. Despite having access to 2048 turns, the agent uses only 80.1 turns on average, indicating adaptive termination once sufficient information is gathered rather than exhaustively consuming the budget. Notably, the growth pattern of average turns mirrors the accuracy curve—both increase rapidly in the $2^{4}$-$2^{7}$ range before plateauing—suggesting that exploration depth naturally aligns with task complexity. This sublinear growth in average turns (compared to exponentially increasing budget) demonstrates that the agent develops increasingly efficient search strategies as more interactions become available, rather than simply extending existing patterns.
@@ 4.5 IterResearch as a Effective Prompting Strategy in Long-Horizon Tasks
Having demonstrated IterResearch’s effectiveness as a trained agent, we investigate whether our iterative paradigm can serve as an effective prompting strategy for long-horizon tasks without any training. We compare with ReAct [yao2023react], the prevailing mono-contextual prompting paradigm, using frontier models o3 [openai_o3_o4_mini] and DeepSeek-V3.1 [dpsk_v31].
@@ 5 Conclusion
In this work, we presented IterResearch, a novel iterative deep-research paradigm that addresses the context suffocation and noise contamination plaguing mono-contextual approaches in long-horizon tasks. By extending the Markov Decision Process to deep research with strategic workspace reconstruction and developing Efficiency-Aware Policy Optimization for effective training, we achieved substantial improvements over existing agents (average +14.5pp across six benchmarks). Furthermore, our experiments reveal three transformative insights: this iterative paradigm enables unprecedented interaction scaling to 2048 interactions with dramatic performance gains (3.5% to 42.5%), serves as an effective prompting strategy that improves frontier models by up to 19.2pp, and induces superior exploration behaviors transferable across different paradigms. These findings establish that iteration with strategic synthesis, rather than accumulation, is fundamental to conquering long-horizon reasoning challenges, providing both a powerful agent architecture and a versatile framework applicable across different models and paradigms.
@@ Appendix A Addtional Related Work
Memory Mechanisms in LLMs. Memory mechanisms have emerged as a critical component for extending LLM capabilities beyond single-turn interactions [du2025rethinking, zhang2025learn]. While early works explored explicit memory architectures with separate storage and retrieval modules [yang2024text], recent approaches have focused on memory management for LLM agents. MemoryLLM [hu2025evaluating] and MEM1 [zhou2025mem1] investigate how agents can learn to synthesize and utilize memory across multi-turn interactions, while Memory-R1 [yan2025memory] employs reinforcement learning to train agents for adaptive memory management. MemAgent [yu2025memagent] and MemOS [li2025memos] further advance this direction by introducing memory operating systems that unify representation, scheduling, and evolution of memories as manageable system resources. However, these memory-centric approaches primarily focus on explicit memory module design or retrieval optimization within fixed context windows, fundamentally differing from our approach. IterResearch naturally integrates memory through the evolving report $\mathcal{M}_{t}$ within our Markovian workspace reconstruction—rather than maintaining separate memory modules or databases, our report serves as a compressed, task-focused memory that is seamlessly updated through the agent’s structured decisions. This design eliminates the overhead of explicit memory management while ensuring that memory evolution is intrinsically aligned with the research trajectory, enabling more efficient and coherent long-horizon exploration.
@@ B.3 Extrapolation Beyond Training Horizon
  • Markovian Workspace: Each round’s decision depends only on the current reconstructed state $(q,\mathcal{M}_{t},\{a_{t-1},\text{TR}_{t-1}\})$, not on absolute position $t$ or the full trajectory history. This position-agnostic design ensures that the agent’s decision-making process remains consistent whether at round 10 or round 1000.
@@ Appendix C More Implementation Details
In this section, we provide a comprehensive implementation details of our proposed method. For additional insights and more intricate details, we refer the reader to our Github Repo.
@@ C.1 Algorithmic Framework
9:   $\text{TR}_{t}\leftarrow\mathcal{E}(a_{t})$ $\triangleright$ Execute tool and get response
10:   $s_{t+1}\leftarrow(q,\mathcal{M}_{t+1},\{a_{t},\text{TR}_{t}\})$ $\triangleright$ Reconstruct workspace
11:   $t\leftarrow t+1$
12: return final answer
@@ C.2 Tool Environment
  • Visit (Web Browser): Enables detailed content extraction from specific URLs with goal-oriented summarization. The agent specifies both the target URLs and a specific goal (e.g., "find the methodology section" or "extract statistical results"), allowing focused information extraction. The tool handles both HTML webpages and PDF documents, automatically detecting and parsing the appropriate format. Our summarization model (Qwen3-30B-A3B) processes the raw content with the agent’s goal to produce concise, relevant summaries. We employ Jina Reader [^2] https://jina.ai/ for robust web content extraction.
@@ C.3 Implementation Details
Chat template Qwen [yang2025qwen3]
@@
Since existing LLMs lack inherent capabilities for our iterative deep-research paradigm, we conduct a two-stage data preparation process: Stage 1: High-quality QA Collection. We curate 30K high-quality question-answer pairs from recent web research datasets [li2025websailor, tao2025webshaper, chen2025expanding, webresearcher]. These pairs are filtered based on answer quality, factual accuracy, and research complexity to ensure they require genuine multi-step investigation. Stage 2: Trajectory Synthesis. To bridge the gap between standard QA pairs and our iterative paradigm, we employ Qwen3-235B-A22B [yang2025qwen3] to synthesize research trajectories following our framework. This process yields 110K training trajectories with an average of 3.7 rounds per trajectory, providing rich supervision for learning the iterative research pattern. We utilize Slime [^4] https://github.com/THUDM/slime as our training framework for the initial supervised fine-tuning phase. The detailed hyper-parameters for this phase are presented in Table 4.
@@
Base model Qwen3-30B-A3B [yang2025qwen3]
@@
We employ a strategic data selection process to identify questions with optimal learning potential: (1) Difficulty Calibration: Using the best checkpoint from SFT, we evaluate each of the 30K questions with 5 independent trials, recording success rates. (2) Learning Zone Selection: We retain questions with success rates between 20%-60% (1-3 correct out of 5 attempts), identifying 4,096 questions that fall within the model’s "zone of proximal development"—challenging enough to provide learning signal but achievable enough to generate successful trajectories. Questions that are too easy ($60\%$ success) provide weak learning signals, while overly difficult questions ($20\%$ success) lead to sparse rewards and unstable training. Table 5 summarizes the key hyperparameters used during the reinforcement learning phase. We also use Slime as our RL frameowrk due to its efficient and easy to use.
@@
GAIA [mialon2023gaia] 32
HLE [hle] 64
BrowseComp-zh [bc_zh] 64
BrowseComp [bc_en] 256
@@
We employ LLM-as-judge evaluation following established practices [chen2025cpo]. Specifically, we use Qwen3-235B-A22B to assess answer correctness:

### v2

@@ 3.2.1 Discounted Reward Shaping for Efficiency
While the iterative workspace reconstruction ensures scalable exploration, a critical question remains: how can we train agents to not just explore deeply, but to do so efficiently? We now address this challenge by employing an efficiency-aware training strategy.
In deep research tasks, the agent receives a binary reward signal $R_{T}\in\{0,1\}$ only upon termination, where $R_{T}=1$ if the final answer is correct and 0 otherwise. This terminal-only reward stems from the inherent difficulty of evaluating intermediate research steps—it is challenging to determine the value of any particular search query or exploratory action (Chen et al., 2025a).
However, this sparse signal alone is insufficient for guiding efficient learning, as it treats all successful trajectories equally regardless of their computational cost. An agent that arrives at the correct answer in 5 well-chosen steps should be preferred over one that requires 20 steps of meandering exploration, even if both ultimately succeed. This efficiency consideration is not merely about computational resources: in real-world deployment, each interaction incurs API costs, and unnecessary exploration can lead to increased latency. To address these issues, we adapt the standard episodic reward formulation from Sutton and Barto (2018). By utilizing geometric discounting, we create an implicit efficiency pressure:
@@ 3.2.2 Policy Optimization with Multi-Round Trajectories
This approach ensures minimal data loss (typically $1\%$ of samples) while maintaining uniform sampling across trajectories. To optimize IterResearch, we combine these efficiency-aware rewards and adaptive downsampling into a unified training strategy (EAPO), implemented on top of the Group Sequence Policy Optimization (GSPO) algorithm (Zheng et al., 2025a), enabling stable training on variable-length trajectories:
@@
where all $\sum_{i=1}^{G}T_{i}$ rounds from the $G$ trajectories for question $q$ form one group, with normalized advantages computed across all samples within this group $\hat{A}_{i,t}=\frac{r_{i,t}-\mu_{r}}{\sigma_{r}}$, $\mathcal{Q}$ is the training set, and $\rho_{i,t}(\theta)$ is the importance ratio based on sequence likelihood (Zheng et al., 2023).
@@ 4.1 Experimental Setup
Datasets. To rigorously assess the effectiveness of our IterResearch, we evaluate on six challenging benchmarks including Humanity’s Last Exam (HLE) (Phan et al., 2025), BrowseComp (Wei et al., 2025a), BrowseComp-zh (Zhou et al., 2025a), GAIA (Mialon et al., 2023), Xbench-DeepSearch (Xbench-Team, 2025), SEAL-0 (Pham et al., 2025). These benchmarks comprehensively assess the essential capabilities for effective deep research in multi-step tool use, web navigation, complex reasoning, long-horizon information-seeking, and cross-lingual synthesis.
Baselines. We comprehensively compare our IterResearch against state-of-the-art methods including: (1) Direct Inference: We evaluate frontier LLMs including GPT-4o and GPT-4.1 (Hurst et al., 2024), o4-mini (OpenAI, 2025b), and DeepSeek-R1-0528 (Guo et al., 2025). (2) Proprietary Deep-Research System: We compare with commercial deep-research systems including OpenAI’s Deep Research (OpenAI, 2025a), Perplexity Research (Perplexity, 2025), Gemini Deep Research (Google, 2025), Grok3-ResearchSearch (xAI, 2025), and Kimi-Researcher (MoonshotAI, 2025). (3) Open-source Agents: Recent open-source deep-research agents including Search-o1 (Li et al., 2025b), WebThinker (Li et al., 2025c), WebDancer (Wu et al., 2025), WebSailor (Li et al., 2025a), Asearcher (Gao et al., 2025), and MiroThinker (MiroMindAI, 2025).
Implementation Details. We implement our IterResearch using Qwen3-30B-A3B (Yang et al., 2025) as the backbone model, considering both model performance and computational efficiency. Our training follows a two-stage process: we first employ rejection sampling fine-tuning (RFT) (Yuan et al., 2023) to equip the model with our iterative deep-research paradigm capabilities, then apply reinforcement learning to further enhance its search strategy and reasoning abilities. For brevity, we provide comprehensive training details and hyperparameters in Appendix C.3.
@@ 4.4 Scaling on Interaction
First, performance scales gracefully with interaction budget. Accuracy improves from 5.5% with only 2 turns to 50.1% at 2048 turns, with the steepest gains occurring between $2^{4}$ and $2^{7}$ turns. This demonstrates that complex information-seeking tasks genuinely benefit from extended exploration—a capability that mono-contextual approaches cannot provide due to inevitable context overflow. Notably, 2048 turns represents an extreme challenge that is currently infeasible for mono-contextual agents due to catastrophic context accumulation, yet our approach operates smoothly within its constant 40K token workspace through iterative state reconstruction. Second, the agent learns intelligent resource allocation. Despite having access to 2048 turns, the agent uses only 80.1 turns on average, indicating adaptive termination once sufficient information is gathered rather than exhaustively consuming the budget. Notably, the growth pattern of average turns mirrors the accuracy curve—both increase rapidly in the $2^{4}$-$2^{7}$ range before plateauing—suggesting that exploration depth naturally aligns with task complexity. This sublinear growth in average turns (compared to exponentially increasing budget) demonstrates that the agent develops increasingly efficient search strategies as more interactions become available, rather than simply extending existing patterns.
@@ 4.5 IterResearch as a Effective Prompting Strategy in Long-Horizon Tasks
Having demonstrated IterResearch’s effectiveness as a trained agent, we investigate whether our iterative paradigm can serve as an effective prompting strategy for long-horizon tasks without any training. We compare with ReAct (Yao et al., 2023), the prevailing mono-contextual prompting paradigm, using frontier models o3 (OpenAI, 2025b) and DeepSeek-V3.1 (DeepSeek, 2025).
@@ 5 Conclusion
In this work, we presented IterResearch, a novel iterative deep-research paradigm that addresses the context suffocation and noise contamination plaguing mono-contextual approaches in long-horizon tasks. By iterative workspace reconstruction and developing Efficiency-Aware Policy Optimization for effective training, we achieved substantial improvements over existing agents (average +14.5pp across six benchmarks). Furthermore, our experiments reveal three transformative insights: this iterative paradigm enables unprecedented interaction scaling to 2048 interactions with dramatic performance gains (3.5% to 42.5%), serves as an effective prompting strategy that improves frontier models by up to 19.2pp, and induces superior exploration behaviors transferable across different paradigms. These findings establish that iteration with strategic synthesis, rather than accumulation, is fundamental to conquering long-horizon reasoning challenges, providing both a powerful agent architecture and a versatile framework applicable across different models and paradigms.
Acknowledgments
This work was supported by Alibaba Research Intern Program. This paper was partially supported by the National Natural Science Foundation of China No. 92470205 and Beijing Major Science and Technology Project under Contract No. Z251100008425002.
@@ Appendix A Addtional Related Work
Memory Mechanisms in LLMs. Memory mechanisms have emerged as a critical component for extending LLM capabilities beyond single-turn interactions (Du et al., 2025; Zhang et al., 2025; Chen et al., 2025b). While early works explored explicit memory architectures with separate storage and retrieval modules (Yang et al., 2024), recent approaches have focused on memory management for LLM agents. MemoryLLM (Hu et al., 2025) and MEM1 (Zhou et al., 2025b) investigate how agents can learn to synthesize and utilize memory across multi-turn interactions, while Memory-R1 (Yan et al., 2025) employs reinforcement learning to train agents for adaptive memory management. MemAgent (Yu et al., 2025) and MemOS (Li et al., 2025d) further advance this direction by introducing memory operating systems that unify representation, scheduling, and evolution of memories as manageable system resources. However, these memory-centric approaches primarily focus on explicit memory module design or retrieval optimization within fixed context windows, fundamentally differing from our approach. IterResearch naturally integrates memory through the evolving report $\mathcal{M}_{t}$ within our Markovian workspace reconstruction—rather than maintaining separate memory modules or databases, our report serves as a compressed, task-focused memory that is seamlessly updated through the agent’s structured decisions. This design eliminates the overhead of explicit memory management while ensuring that memory evolution is intrinsically aligned with the research trajectory, enabling more efficient and coherent long-horizon exploration.
Iterative Reasoning and Convergent Paradigms. Concurrently, similar iterative paradigms have emerged in different domains, most notably in algorithmic discovery with systems like AlphaEvolve (Novikov et al., 2025). AlphaEvolve, designed for coding tasks, shares a strong conceptual architecture with IterResearch: it maintains a main objective (the task), an “evolving report” of high-level ideas (analogous to our $\mathcal{M}_{t}$), and an immediate context (the code). Despite this architectural parallel, the domains are fundamentally different. AlphaEvolve operates on structured code and receives verification from unit tests, whereas IterResearch is designed for the unstructured, noisy, and dynamic environment of web research. We believe this connection strengthens our core thesis. The ”convergent evolution” of two distinct, long-horizon domains (algorithmic discovery and deep web research) independently arriving at an iterative synthesis paradigm with an evolving report strongly suggests this is a fundamental and highly effective solution to the limitations of mono-contextual reasoning. It indicates that our IterResearch paradigm is not a niche method, but a general and advanced architecture for complex, long-horizon AI.
@@ B.3 Extrapolation Beyond Training Horizon
  • Workspace: Each round’s decision depends only on the current reconstructed state $(q,\mathcal{M}_{t},\{a_{t-1},\text{TR}_{t-1}\})$, not on absolute position $t$ or the full trajectory history. This position-agnostic design ensures that the agent’s decision-making process remains consistent whether at round 10 or round 1000.
@@ Appendix C More Implementation Details
In this section, we provide a comprehensive implementation details of our proposed method. For additional insights and more intricate details, we refer the reader to our supplementary materials.
@@ C.1 Algorithmic Framework
9:  end if
10:  $\text{TR}_{t}\leftarrow{\mathcal{E}}(a_{t})$ $\triangleright$ Execute tool and get response
11:  $s_{t+1}\leftarrow(q,{\mathcal{M}}_{t+1},\{a_{t},\text{TR}_{t}\})$ $\triangleright$ Reconstruct workspace
12:  $t\leftarrow t+1$
13:end while
14:return final answer
@@ C.2 Tool Environment
  • Visit (Web Browser): Enables detailed content extraction from specific URLs with goal-oriented summarization. The agent specifies both the target URLs and a specific goal (e.g., ”find the methodology section” or ”extract statistical results”), allowing focused information extraction. The tool handles both HTML webpages and PDF documents, automatically detecting and parsing the appropriate format. Our summarization model (Qwen3-30B-A3B) processes the raw content with the agent’s goal to produce concise, relevant summaries. We employ Jina Reader [^2] https://jina.ai/ for robust web content extraction.
@@ C.3 Implementation Details
Chat template Qwen (Yang et al., 2025)
@@
Since existing LLMs lack inherent capabilities for our iterative deep-research paradigm, we conduct a two-stage data preparation process: Stage 1: High-quality QA Collection. We curate 30K high-quality question-answer pairs from recent web research datasets (Li et al., 2025a; Tao et al., 2025; Chen et al., 2025c; Qiao et al., 2025). These pairs are filtered based on answer quality, factual accuracy, and research complexity to ensure they require genuine multi-step investigation. Stage 2: Trajectory Synthesis. To bridge the gap between standard QA pairs and our iterative paradigm, we employ Qwen3-235B-A22B (Yang et al., 2025) to synthesize research trajectories following our framework. This process yields 110K training trajectories with an average of 3.7 rounds per trajectory, providing rich supervision for learning the iterative research pattern. We utilize Slime [^4] https://github.com/THUDM/slime as our training framework for the initial supervised fine-tuning phase. The detailed hyper-parameters for this phase are presented in Table 4.
@@
Base model Qwen3-30B-A3B (Yang et al., 2025)
@@
We employ a strategic data selection process to identify questions with optimal learning potential: (1) Difficulty Calibration: Using the best checkpoint from SFT, we evaluate each of the 30K questions with 5 independent trials, recording success rates. (2) Learning Zone Selection: We retain questions with success rates between 20%-60% (1-3 correct out of 5 attempts), identifying 4,096 questions that fall within the model’s ”zone of proximal development”—challenging enough to provide learning signal but achievable enough to generate successful trajectories. Questions that are too easy ($60\%$ success) provide weak learning signals, while overly difficult questions ($20\%$ success) lead to sparse rewards and unstable training. Table 5 summarizes the key hyperparameters used during the reinforcement learning phase. We also use Slime as our RL frameowrk due to its efficient and easy to use.
@@
GAIA (Mialon et al., 2023) 32
HLE (Phan et al., 2025) 64
BrowseComp-zh (Zhou et al., 2025a) 64
BrowseComp (Wei et al., 2025a) 256
@@
We employ LLM-as-judge evaluation following established practices (Chen et al., 2025a). Specifically, we use Qwen3-235B-A22B to assess answer correctness:
@@
Baselines.
The baseline performance figures for prompting-based comparisons in Table 1 (e.g., ReAct on BrowseComp) were sourced directly from the State-of-the-Art (SOTA) results reported in their original papers. We adopted this methodology for two critical reasons: (1) Many of these prior SOTA works do not publicly release their exact optimized prompts. (2) Therefore, directly reporting their published score is the most rigorous and fair method for comparison. This ensures our results are benchmarked against the true SOTA, not a potentially weaker or un-optimized re-implementation.
C.4 Average Token Budgets
To provide further insight into the reproducibility and computational cost of our method, we detail the average generation tokens required for the IterResearch-30B-A3B agent to complete tasks across the six primary benchmarks.
It is important to note that this data **excludes** tokens returned by tools (e.g., the content of web pages or search results). It represents only the tokens consumed by the agent’s own ”Think” and ”Report” steps. This figure thus reflects the actual computational cost of the agent’s internal reasoning.
Table 7: Average generation tokens (excluding tool responses) for IterResearch-30B-A3B across benchmarks.
Benchmark HLE BC BC-zh GAIA Xbench-DS Seal-0
Avg. Tokens 31,426 376,349 81,136 33,198 31,600 28,186
As shown in Table 7, the required reasoning cost (i.e., generation tokens) varies dramatically across benchmarks.
  • BrowseComp (BC) consumes a token budget (approx. 376k) that is an order of magnitude larger than all other benchmarks. This provides strong empirical evidence of its extreme complexity as a long-horizon task, requiring the agent to perform extensive iteration, thinking, and report synthesis to succeed in the complex web environment. This aligns with our scaling experiments (Figure 3), where performance continually improves as more interaction steps (and thus more generation tokens) are permitted.
  • BrowseComp-zh (BC-zh), at approx. 81k tokens, is the second most complex task.
  • The other four benchmarks (GAIA, HLE, Seal-0, Xbench-DS) are of a similar magnitude (approx. 28k-33k tokens), suggesting that while they are complex multi-step reasoning tasks, they require significantly less iterative synthesis compared to the long-horizon ‘BC‘ benchmarks.
This analysis demonstrates that our method’s performance gains (e.g., in Table 2) do not stem from uncontrolled token expenditure. Instead, they are achieved by a superior paradigm (IterResearch) operating within a manageable, task-dependent reasoning budget.